Consider a uniport system where a carrier protein
Answer #1

\DeltaG0   = 11.1Kj/mole = 11100J/mole

T      = 25+ 273 = 298K

A outside --------> A inside

K   = [A] inside/[A] outside

\DeltaG0    = -RTlnK

-11100 = -8.314*298*2.303log [A] inside/[A] outside

log [A] inside/[A] outside    = 11100/-5705.8483

log [A] inside/[A] outside     = -1.945

[A] inside/[A] outside           =-10^1.945

[A] inside/[A] outside         = -88.1 >>>> answer

b) option b is correct

gelta G is negative reaction is spontaneous